# R-enum Revisited: Speedup and Extension for Context-Sensitive Repeats and Net Frequencies

### v1

@@ 6 Conclusions
We improved the time complexity of r-enum, and extended its enumeration power for other context-sensitive repeats in the literature. We also proposed the first data structure of size $O(r)$ to support NF-queries in optimal time. As a combinatorial property, we formally proved that the total number of net occurrences, as well as MUSs, is upper bounded by $O(r)$. We finally remark that our work leads to the first algorithm to compute the sorted list of net occurrences and MUSs in $O(n)$ time and $O(r)$ space: We enumerate all net occurrences by Theorem˜13, and sort them in $O(n)$ time and $O(r)$ space by [25, Lemma 13].
Acknowledgements
Tomohiro I was supported by JSPS KAKENHI (Grant Numbers JP24K02899) and JST AIP Acceleration Research JPMJCR24U4, Japan.

### v2

@@ 6 Conclusions
We improved the time complexity of r-enum, and extended its enumeration power for other context-sensitive repeats in the literature. We also proposed the first data structure of size $O(r)$ to support NF-queries in optimal time, which can be regarded as an $O(r)$-space representation of $\mathcal{S}^{\mathsf{nsmr}}$ with the help of RLBWTs. As a combinatorial property, we formally proved that the total number of net occurrences, as well as MUSs, is upper bounded by $O(r)$.
We note that our enumeration algorithm runs independently of the alphabet size $\sigma$ thanks to the output-optimal range distinct queries of Lemma˜2.3. The space usage depending on $\sigma$ is hidden behind $O(r)$. We removed the dependence on $\sigma$ in the NF-query time by using a dictionary with worst-case constant-time lookup.
Finally, we remark that our work leads to the first algorithm to compute the sorted list of net occurrences and MUSs in $O(n)$ time and $O(r)$ space: We enumerate all net occurrences by Theorem˜4.2, and sort them in $O(n)$ time and $O(r)$ space by [2020NishimotoT_FasterQueriesOnBwtRuns_X3, Lemma 13].